### v2

@@ 5.3 Other Domain-Specific 3D Applications
QINet [168] applies RL to 3D point cloud completion using an actor-critic framework. The agent interacts with a latent-space GAN, refining noise vectors to reconstruct complete shapes from corrupted inputs. Rewards combine geometric accuracy (IoU) and latent consistency, enabling robust completion even with large missing regions. RLSS [169] addresses sequential 3D scene generation. It formulates scene construction as an RL problem where states encode 2D layouts and object placements, and actions select and position objects. Using PPO, the agent is trained with rewards that enforce physical constraints and design objectives. The method enables diverse and realistic indoor layouts with minimal supervision. Zhao et al. [170] propose a scene-aware RL framework for generating human motion in 3D environments. PPO is used to train policies over a latent motion space. States capture geometry, body pose, and interaction goals, while rewards encourage goal completion, contact accuracy, and collision avoidance. The approach supports complex interactions and fine-grained path control. Bailando [171] introduces a transformer-based RL framework for 3D dance generation. A choreographic memory encodes quantized poses, and a GPT-style actor-critic predicts future motion. Beat-alignment rewards ensure synchronization with music and consistency between body parts, producing realistic and stylistically coherent dances. Text2Stereo [172] tackles text-to-stereo image generation. It fine-tunes Stable Diffusion on vertically stacked stereo pairs and uses RL via AlignProp. Differentiable rewards guide optimization toward stereo consistency and prompt alignment, with human preferences integrated into the feedback signal.
@@
Object Generation with Design Constraint [155] 2020 RE Constraint-aware 3D Objects Generation ML Neighboring voxels (6 or 26 directions) moving or specific LEGO blocks placing, depending on the experiment +1 (inside geometry), -0.5/-1 (self-intersection / outside) Local voxel window (e.g., 5x5x5) around the agent Generates topologically consistent, stylistically varied 3D objects and LEGO assemblies, topologically consistent designs meeting constraints
@@
DeepMesh [156] 2025 Arxiv Mesh Generation DPO Vertex/face token prediction in hierarchical blocks Human preference + geometric metrics (Chamfer Distance) Local voxel window (e.g., 5×5×5) + patch connectivity. SOTA in geometric fidelity, visual appeal, generation efficien and face count (30k+)
@@
Mesh-RFT [157] 2025 Arxiv Mesh Generation Masked DPO Face-level token prediction with quality-aware masking BER (Boundary Edge Ratio) + TS (Topology Score) + HD (Hausdorff Distance) metrics Tokenized mesh sequences + local quality masks. SOTA in both geometric accuracy and visual fidelity
@@
Point Cloud Scene Completion from Single RGB-D Image [158] 2023 TPAMI Colored Semantic Point Cloud Scene Completion A3C Next-best-view selection Accuracy, hole-filling, and point cloud recovery Updated point cloud at each iteration SOTA in scene completion and scene segmentation accuracy
@@
3D Shapes Modeling [10] 2020 ECCV 3D Shape Modeling IL + RL Primitive manipulation and mesh editing IoU-based and parsimony rewards Shape reference, primitives/edge loops, step encoding Structurally meaningful and regular meshes
@@
DreamReward [9] 2024 ECCV Text-to-3d Generation RLHF Optimizing 3D representations using human feedback Reward3D (human preference scoring) 3D representations SOTA in high-fidelity, human-aligned 3D generation
@@
DreamDPO [160] 2025 Arxiv Text-to-3D Generation DPO Optimizing 3D representations via pairwise comparisons Preference scores from reward/LMM models 3D representations SOTA in high-quality, controllable 3D generation
@@
DreamCS [161] 2025 Arxiv Text-to-3D Generation CS divergence Optimizing 3D representations via unpaired preference learning RewardCS (3D geometry-aware scoring) 3D representations SOTA in high-fidelity, geometrically consistent 3D generation
@@
Nabla-R2D3 [163] 2025 Arxiv Text-to-3D Generation Gradient-informed RL Finetune Sampling and rendering multi-view images 2D appearance and geometry scores. Latent 3D shape representations Propose a method Nabla-R2D3 that can effectively, efficiently and robustly finetune 3D-native generative models from 2D reward models with better preference alignment, better text-object alignment and fewer geometric artifacts
@@
Carve3d [164] 2024 CVPR Text-to-Multiview Generation On-policy Policy Gradient, KL regularization Denoising steps in diffusion process MRC (Multi-view Reconstruction Consistency) metric Text prompt, timestep, noisy image Superior multi-view consistency and NeTF reconstruction quality
@@
HFDream [165] 2024 OpenReview Text-to-Multiview Generation LoRA, SFT Generating view-aligned images Human-annotated view alignment score Text prompt and diffusion model output Improved 3D consistency and text alignment
@@
MVReward [166] 2025 AAAI Image-to-Multiview Generation MVP Denoising steps in diffusion process Human preference score (MVReward) Image prompt and noisy multi-view images Improved human-aligned 3D generation quality
@@
Refining Few-Step Text-to-Multiview Diffusion [167] 2025 Arxiv Text-to-Multiview Generation MV-ZigAL, Constrained Optimization Denoising steps in diffusion process Joint-view scores including alignment quality, geometry quality, texture quality, and overall quality, and single-view fidelity scores Multiview-aware MDP (Markov decision process) with text prompt and viewpoint embeddings Improved fidelity and consistency while preserving few-step T2MV diffusion baseline efficiency
@@
QINet [168] 2022 TGRS Point Cloud Completion DDPG Revised noise of z (clean latent code) Weighted sum of reconstruction and latent rewards Incomplete representation in latent space (IRLS) SOTA in shape completeness, perceptual quality, shape resolution and category distinction
@@
RLSS [169] 2022 WACV 3D Indoor Scene Generation with Predefined Design Objects PPO Object placement and selection Constraint-driven rewards that include multiple conditions such as successful condition, count of objects, and failure conditions 2D scene representation with object metadata High-quality, diverse scene generation with fast synthesis
@@
Synthesizing diverse human motions in 3d indoor scenes [170] 2023 ICCV Human Motions Synthesis PPO Latent variables of CVAE-based generative motion primitive Goal-reaching, foot-ground contact, and penetration avoidance configurations of 3D scene geometry, virtual human body, and intended goals to interact with Realistic and diverse human-scene interactions with low penetration, high foot-ground contact accuracy, and efficient task completion times, outperforming baselines in perceptual studies
@@
Bailando [171] 2022 CVPR 3D Dance Generation Actor-Critic Finetune Predicting next dance pose Beat-align reward and half-body consistency reward A sequence of dance poses SOTA in motion quality, diversity, and music-dance alignment
@@ 6.3 Future Work
Abbreviations
  • RL: Reinforcement Learning
  • MDP: Markov Decision Process
  • PPO: Proximal Policy Optimization
  • GRPO: Group Relative Policy Optimization
  • DPO: Direct Preference Optimization
  • CFG: Classifier-Free Guidance
  • RLHF: Reinforcement Learning with Human Feedback
  • VPO: Visual Preference Optimization
  • IPO: Iterative Preference Optimization
  • NPO: Negative Preference Optimization
  • LPIPS: Learned Perceptual Image Patch Similarity
  • SSIM: Structural Similarity Index
  • GAN: Generative Adversarial Network
  • NeRF: Neural Radiance Fields
  • GPT: Generative Pre-trained Transformer
  • CLIP: Contrastive Language–Image Pre-training
  • BLIP-2: Bootstrapped Language–Image Pretraining v2
  • YOLO: You Only Look Once
  • A3C: Asynchronous Advantage Actor-Critic
  • IoU: Intersection over Union
  • RGB-D: Red–Green–Blue with Depth
  • BER: Boundary Edge Ratio
  • TS: Topology Score
  • MV: Multi-View
  • RLFT: Reinforcement Learning Fine-Tuning
  • MRC: Multi-view Reconstruction Consistency
  • ZMV: Zigzag Multi-View (in ZMV-Sampling)
  • MV-ZigAL: Multi-View Zigzag Advantage Learning
  • DDT: Diffusion Timestep Tokenizer
  • LRM: Large Reconstruction Model
Acknowledgements
The author would like to thank colleagues for their valuable feedback during the preparation of this survey.

### v3

@@ 5.3 Other Domain-Specific 3D Applications
QINet [zhang2022qinet] applies RL to 3D point cloud completion using an actor-critic framework. The agent interacts with a latent-space GAN, refining noise vectors to reconstruct complete shapes from corrupted inputs. Rewards combine geometric accuracy (IoU) and latent consistency, enabling robust completion even with large missing regions. RLSS [ostonov2022rlss] addresses sequential 3D scene generation. It formulates scene construction as an RL problem where states encode 2D layouts and object placements, and actions select and position objects. Using PPO, the agent is trained with rewards that enforce physical constraints and design objectives. The method enables diverse and realistic indoor layouts with minimal supervision. Zhao et al. [zhao2023synthesizing] propose a scene-aware RL framework for generating human motion in 3D environments. PPO is used to train policies over a latent motion space. States capture geometry, body pose, and interaction goals, while rewards encourage goal completion, contact accuracy, and collision avoidance. The approach supports complex interactions and fine-grained path control. Bailando [siyao2022bailando] introduces a transformer-based RL framework for 3D dance generation. A choreographic memory encodes quantized poses, and a GPT-style actor-critic predicts future motion. Beat-alignment rewards ensure synchronization with music and consistency between body parts, producing realistic and stylistically coherent dances. Text2Stereo [garg2025text2stereo] tackles text-to-stereo image generation. It fine-tunes Stable Diffusion on vertically stacked stereo pairs and uses RL via AlignProp. Differentiable rewards guide optimization toward stereo consistency and prompt alignment, with human preferences integrated into the feedback signal.
@@
Object Generation with Design Constraint [akizuki2020generative] 2020 RE Constraint-aware 3D Objects Generation ML Neighboring voxels (6 or 26 directions) moving or specific LEGO blocks placing, depending on the experiment +1 (inside geometry), -0.5/-1 (self-intersection / outside) Local voxel window (e.g., 5x5x5) around the agent Generates topologically consistent, stylistically varied 3D objects and LEGO assemblies, topologically consistent designs meeting constraints
@@
DeepMesh [zhao2025deepmesh] 2025 Arxiv Mesh Generation DPO Vertex/face token prediction in hierarchical blocks Human preference + geometric metrics (Chamfer Distance) Local voxel window (e.g., 5×5×5) + patch connectivity. SOTA in geometric fidelity, visual appeal, generation efficien and face count (30k+)
@@
Mesh-RFT [liu2025mesh] 2025 Arxiv Mesh Generation Masked DPO Face-level token prediction with quality-aware masking BER (Boundary Edge Ratio) + TS (Topology Score) + HD (Hausdorff Distance) metrics Tokenized mesh sequences + local quality masks. SOTA in both geometric accuracy and visual fidelity
@@
Point Cloud Scene Completion from Single RGB-D Image [zhang2023point] 2023 TPAMI Colored Semantic Point Cloud Scene Completion A3C Next-best-view selection Accuracy, hole-filling, and point cloud recovery Updated point cloud at each iteration SOTA in scene completion and scene segmentation accuracy
@@
3D Shapes Modeling [lin2020modeling] 2020 ECCV 3D Shape Modeling IL + RL Primitive manipulation and mesh editing IoU-based and parsimony rewards Shape reference, primitives/edge loops, step encoding Structurally meaningful and regular meshes
@@
DreamReward [ye2024dreamreward] 2024 ECCV Text-to-3d Generation RLHF Optimizing 3D representations using human feedback Reward3D (human preference scoring) 3D representations SOTA in high-fidelity, human-aligned 3D generation
@@
DreamDPO [zhou2025dreamdpo] 2025 Arxiv Text-to-3D Generation DPO Optimizing 3D representations via pairwise comparisons Preference scores from reward/LMM models 3D representations SOTA in high-quality, controllable 3D generation
@@
DreamCS [zou2025dreamcs] 2025 Arxiv Text-to-3D Generation CS divergence Optimizing 3D representations via unpaired preference learning RewardCS (3D geometry-aware scoring) 3D representations SOTA in high-fidelity, geometrically consistent 3D generation
@@
Nabla-R2D3 [liu2025nabla] 2025 Arxiv Text-to-3D Generation Gradient-informed RL Finetune Sampling and rendering multi-view images 2D appearance and geometry scores. Latent 3D shape representations Propose a method Nabla-R2D3 that can effectively, efficiently and robustly finetune 3D-native generative models from 2D reward models with better preference alignment, better text-object alignment and fewer geometric artifacts
@@
Carve3d [xie2024carve3d] 2024 CVPR Text-to-Multiview Generation On-policy Policy Gradient, KL regularization Denoising steps in diffusion process MRC (Multi-view Reconstruction Consistency) metric Text prompt, timestep, noisy image Superior multi-view consistency and NeTF reconstruction quality
@@
HFDream [choi2024hfdream] 2024 OpenReview Text-to-Multiview Generation LoRA, SFT Generating view-aligned images Human-annotated view alignment score Text prompt and diffusion model output Improved 3D consistency and text alignment
@@
MVReward [wang2025mvreward] 2025 AAAI Image-to-Multiview Generation MVP Denoising steps in diffusion process Human preference score (MVReward) Image prompt and noisy multi-view images Improved human-aligned 3D generation quality
@@
Refining Few-Step Text-to-Multiview Diffusion [zhang2025refining] 2025 Arxiv Text-to-Multiview Generation MV-ZigAL, Constrained Optimization Denoising steps in diffusion process Joint-view scores including alignment quality, geometry quality, texture quality, and overall quality, and single-view fidelity scores Multiview-aware MDP (Markov decision process) with text prompt and viewpoint embeddings Improved fidelity and consistency while preserving few-step T2MV diffusion baseline efficiency
@@
QINet [zhang2022qinet] 2022 TGRS Point Cloud Completion DDPG Revised noise of z (clean latent code) Weighted sum of reconstruction and latent rewards Incomplete representation in latent space (IRLS) SOTA in shape completeness, perceptual quality, shape resolution and category distinction
@@
RLSS [ostonov2022rlss] 2022 WACV 3D Indoor Scene Generation with Predefined Design Objects PPO Object placement and selection Constraint-driven rewards that include multiple conditions such as successful condition, count of objects, and failure conditions 2D scene representation with object metadata High-quality, diverse scene generation with fast synthesis
@@
Synthesizing diverse human motions in 3d indoor scenes [zhao2023synthesizing] 2023 ICCV Human Motions Synthesis PPO Latent variables of CVAE-based generative motion primitive Goal-reaching, foot-ground contact, and penetration avoidance configurations of 3D scene geometry, virtual human body, and intended goals to interact with Realistic and diverse human-scene interactions with low penetration, high foot-ground contact accuracy, and efficient task completion times, outperforming baselines in perceptual studies
@@
Bailando [siyao2022bailando] 2022 CVPR 3D Dance Generation Actor-Critic Finetune Predicting next dance pose Beat-align reward and half-body consistency reward A sequence of dance poses SOTA in motion quality, diversity, and music-dance alignment